# Gold-medalist Performance in Solving Olympiad Geometry with AlphaGeometry2

### v2

@@ Abstract
We present AlphaGeometry2, a significantly improved version of AlphaGeometry introduced in Trinh et al. (2024), which has now surpassed an average gold medalist in solving Olympiad geometry problems. To achieve this, we first extend the original AlphaGeometry language to tackle harder problems involving movements of objects, and problems containing linear equations of angles, ratios, and distances. This, together with support for non-constructive problems, has markedly improved the coverage rate of the AlphaGeometry language on International Math Olympiads (IMO) 2000-2024 geometry problems from 66% to 88%. The search process of AlphaGeometry2 has also been greatly improved through the use of Gemini architecture for better language modeling, and a novel knowledge-sharing mechanism that enables effective communication between search trees. Together with further enhancements to the symbolic engine and synthetic data generation, we have significantly boosted the overall solving rate of AlphaGeometry2 to 84% for all geometry problems over the last 25 years, compared to 54% previously. AlphaGeometry2 was also part of the system that achieved silver-medal standard at IMO 2024 https://dpmd.ai/imo-silver. Last but not least, we report progress towards using AlphaGeometry2 as a part of a fully automated system that reliably solves geometry problems directly from natural language input.
@@ 1 Introduction
Advancing reasoning capabilities of artificial intelligence (AI) systems, especially in tackling complex mathematical problems, has been a decade-long holy grail of AI research. Even with the rapid developments of large language models, most systems until now still struggle with fundamental understanding of geometric objects (e.g., points, lines, circles) and their interactions. This had motivated our previous work, AlphaGeometry Trinh et al. (2024), a neuro-symbolic system capable of solving Euclidean geometry problems at the Olympiad level. Specifically, AlphaGeometry (AG1) combines the power of a neural language model, that is creative in suggesting insights for tackling challenging problems, with a symbolic engine, that uses a formal language to describe geometric constructs and deduction rules to reliably derive new properties. Trained on 100M synthetically generated examples of diverse levels of complexity, AG1 was able to solve geometry problems at the level of the International Mathematical Olympiad (IMO), a prestigious competition for the world’s brightest high-school mathematicians.
@@
Broadly speaking, there are two main approaches for automatically solving geometry problems. The first group is about bashing the geometry problems algebraically with Wu’s method Chou (1985); Wu (2008), Area method Chou et al. (1993, 1994), or Gröbner bases Kapur (1986a, b). The second group, on the other hand, relies on synthetic techniques, such as Deduction database Chou et al. (2000), or Full angle method Chou et al. (1996). AG1 uses the latter, as a more human-like approach suitable for transferring the research knowledge to other domains. The neuro-symbolic system that AG1 employed has demonstrated a significant step towards mastering the domain of Euclidean geometry. However, despite its success, AG1 exhibited limitations in several key areas. Its performance was constrained by the scope of its domain-specific language, the efficiency of the symbolic engine, and the capacity of the initial language model. As a result, when considering all the recent IMO geometry problems from the year 2000 until now, AG1 can only achieve a solving rate of 54%.
@@
This paper introduces AlphaGeometry2 (AG2), a substantial upgrade that addresses the aforementioned limitations and significantly enhances performance. First, we expand the domain language to encompass a wider range of geometric concepts, including locus theorems with moving constructs and linear equations of angles, ratios, and distances. We also introduce a significantly faster and more robust symbolic engine, incorporating optimizations such as a reduced rule set and enhanced handling of double points. AG2 leverages a more powerful language model based on Gemini Team Gemini (2024) and trained on an order of magnitude larger and more diverse dataset. To further improve performance, we develop a novel search algorithm, Shared Knowledge Ensemble of Search Trees (SKEST), that explores a broader range of auxiliary construction strategies and employs a knowledge-sharing mechanism to expand and accelerate the search process. Finally, we make progress towards building a fully automated and reliable system that solves geometry problems in natural language. To do this we utilize Gemini to translate problems from natural language into the AlphaGeometry language and implement a new automated diagram generation algorithm.
@@
  • Stronger and faster symbolic engine: introducing an optimized rule set, better handling of double points, and a two-order of magnitude faster implementation in C++.
@@ 2 More general domain language
First introduced in Trinh et al. (2024), AG1 uses a simple domain specific language that consists of nine basic “predicates” listed in Table 1. While these predicates are sufficient to cover 66% of all geometry problems for 2000-2024 IMO, AG1 language does not allow talking about linear equations, movements of points/lines/circles or common problems such as ”Find the angle …“. Below we explain how AG2 addresses these and other challenges.
@@
eqangle a b c d e f g h Directed angle between $AB$ and $CD$ is the same as the one between $EF$ and GHGH
@@
3 Automated problem formalization and diagram generation
Automated formalization.
A major weakness of AlphaGeometry, and other similar neuro-symbolic systems, is the need to manually transform input problems from natural language into a domain specific language. For example, a simple geometry problem in natural language, “Given a triangle $ABC$ with two equal sides $AB=AC$, prove that angles $B$ and $C$ are equal”, becomes triangle a b c; a b = a c ? eqangle b a b c c b c a in the AlphaGeometry domain language.
Automating this process, called formalization, is an active area of research (see for example, Szegedy (2020); Wu et al. (2022); Jiang et al. (2023); Poiroux et al. (2024)). It is a significantly more complicated problem compared to a translation between human languages. While translation aims to preserve meaning, formalization frequently requires re-formulating the original problem into an alternative form, and sometimes disambiguating the nuances in the original problem statement. Automated formalization (auto-formalization), therefore, demands significant background knowledge and problem-solving skills on its own. Given that recently foundation models started demonstrating such capabilities, we use one such model, Gemini Team Gemini (2024), to automate the problem formalization for AlphaGeometry. We start by manually translating several dozens of geometry problems into the AG language. Then we use these examples to write a few-shot prompt asking Gemini to translate a given geometry problem from natural language into the AG language. We query Gemini five times with this prompt followed by another Gemini call asking to combine these results into one final answer. With this approach we are able to formalize 33 out of 44 formalizable IMO 2000-2024 geometry problems. For easier geometry problems, it is very consistent and makes almost no mistakes.
Automated diagram generation.
Another manual part of our pipeline was diagram generation. In AG1, each point is defined by at most two basic predicates recalled in Table 1, the problem is therefore defined constructively and diagrams can be generated automatically. In AG2, we allow one or multiple points being defined simultaneously by an arbitrary number of predicates, allowing us to also cover non-constructive problems. Consider a non-constructive problem statement, “Let $ABC$ be a triangle with incenter $I$, such that $IA=2IB$ …”, here point $I$ is not only defined as an incenter, i.e. the intersection of two internal bisectors, but also defined by a third predicate $IA=2IB$ and there is no general strategy to construct such four points. Since AG2 covers non-constructive problems, diagram construction becomes a non-trivial part of the pipeline and generally requires human intervention. Similar to Krueger et al. (2021), below we propose a new algorithm to automatically generate diagrams given non-constructive problem specifications.
We start by initializing the points. We have three initialization methods which we alternate in different attempts:
  1. 1.
  2. 2.
    Construction in order.
  3. 3.
    Construction in a heuristically chosen order.
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
This three-stage optimization method is similar to the methodology introduced in Krueger et al. (2021). The final stage addresses the practical limitations encountered when tuning the gradient descent optimization in the original method, where achieving a consistently satisfactory error margin proved challenging.
@@ 4.2 Faster algorithm
The DDAR1 algorithm is processing a list of rules, and tries to apply each rule to all combinations of points. This process involves a candidate searching step, whose time complexity is polynomial in the number of points, and a clause matching step, whose time complexity is exponential in the number of clauses per premise. In theory the worst case for searching similar triangle candidates in AG1 is $O(N^{8})$, which is one of the most time consuming steps. The exponential clause matching step is another expensive step. To make the search more efficient, we take all essential rules, and hard-code search for their application, which reduces the number of queries for the AR sub-engine to at most cubic. Furthermore, we discard the explicit rules for angles and distances (e.g. about perpendicular or parallel lines) – all such deductions happen automatically in the AR engine.
@@ 4.3 Faster implementation
While the new algorithm already significantly accelerates DDAR, we make further speed improvements by implementing its core computation (Gaussian Elimination) in C++. The new C++ library, which is exported into Python via pybind11 Jakob et al. (2017), is over 300 times faster than DDAR1. In order to benchmark the speed improvement, we select a set of 25 IMO problems that cannot be solved by DDAR (see Figure 8), and run the test 50 times on a machine with AMD EPYC 7B13 64 core CPU. While on average DDAR1 finishes its computations in $1179.57\pm 8.055$ seconds, DDAR2 is much faster - finishing in $3.44711\pm 0.05476$ seconds [^1] The average running time may vary depending on the machine status at different times..
@@ 5 Better synthetic training data
Supplementing the symbolic engine with a language model was a key to AG1 success, bringing the solve rate from 14 (pure deduction proofs) to 25 out of 30 selected IMO problems Trinh et al. (2024). This language model was trained on a large amount of algorithmically generated synthetic data. In AG2 we use the same procedure.
Similar to AG1, our synthetic data generation method starts by sampling a random diagram, and uses the symbolic engine to deduce all possible facts from it. For each of the deduced facts, a traceback algorithm is used to extract the corresponding premises, auxiliary points, and deduction steps that prove the fact. Our data generation method deliberately avoids the use of human-crafted problems as initial diagram seeds, and strictly starts from random diagrams. This design choice eliminates the risk of data contamination and allows for the exploration of theorem distributions that may extend beyond established human knowledge. This approach contrasts with methods like TongGeometry Zhang et al. (2024), which rely on human expertise and existing problem diagrams to guide and filter data generation. In AG2, we keep using random diagrams as initial seeds and continue to push for better synthetic training data.
@@
We also improved the speed of the data generation algorithm. Recall that in AG1, we first run deduction closure on random diagrams, and “traceback" to obtain the minimal problem and minimal proof that proves each fact in the closure. To obtain the minimal problem in AG1, we have to exhaustively remove different subsets of points from the problem and rerun DDAR to check provability. Such a search can find the subset with the smallest cardinality, however as an exponential search, it is infeasible for a larger number of points. Therefore, we switched to a greedily discarding algorithm shown in Figure 3, which uses just a linear number of checks for whether a set of points suffices to prove the goal. The greedy algorithm is guaranteed to find a minimal set of points with respect to inclusion as long as the check is monotonic (if $A\subseteq B$, then $\texttt{check\_provable}(A)\Rightarrow\texttt{check\_provable}(B)$). In reality, we also require the pruned set to remain closed under construction dependencies (so that we can still run a random construction). If we incorporate this condition into the check_provable predicate, it stops being monotonic. This difficulty can be fixed by processing the points via the algorithm from Figure 3 in a reverse-topological order (first points that do not depend on any other points, and last the initial points of the construction).
@@ 6 Novel search algorithm
In AG1, we use a simple beam search to discover proofs. In AG2, we design a novel search algorithm, in which several differently configured beam searches are executed in parallel and are allowed to help each other through a knowledge sharing mechanism (see Figure 4). To improve the robustness of our system we use multiple different language models for each search tree configuration. We call this search algorithm Shared Knowledge Ensemble of Search Trees (SKEST).
@@
  • "Classic" search tree: the same beam tree search used in AG1, where a language model is asked to produce one auxiliary point at each node.
  • Tree predicting multiple auxiliary points at each node: a language model is allowed to produce as many auxiliary points as it wants at each tree node. Recall that this is possible because our LM is trained to produce full proofs, starting with auxiliary points and followed by the deduction steps [^2] See a more detailed discussion on producing full proofs with a language model alone in Section E. Note that even though we want our models to generate all necessary auxiliary points in one query, in practice we observe the need to call the model multiple times given previously produced auxiliary points. Allowing the model to produce multiple auxiliary points accelerates finding a solution and effectively increases the tree search depth.
  • Tree predicting different types of aux points uniformly. Recall that LM outputs for auxiliary points look like x00 a : cong a b c d (00) coll a e f (01), i.e. “construct point a such that a b = c d and a e f are collinear". Typically to predict aux points we prompt the language model with the first token x00 and let it generate the rest. Here, instead, we prompt the LM with x00 a : cong, x00 a : coll, x00 a : cyclic, x00 a : perp, etc. to force uniform distribution across the first 4 tokens, and then let the LM generate the rest.
@@ 7.1 Training setup
AG1 language model, a custom transformer, was trained in the unsupervised fashion in two phases: training on problems with and without auxiliary constructions followed by training on only problems that contain auxiliary constructions. For AG2 we leverage the Gemini training pipeline and simplify training to just one phase: unsupervised learning on all data. Our new language model is a sparse mixture-of-expert Transformer-based model that builds on Gemini Team Gemini (2024) and trained on AG2 data described in Section 5. We train multiple models of different size using three training setups:
@@
    Training from scratch with a custom tokenizer in the domain specific language (AG1 setup).
@@
    Fine-tuning already pre-trained custom math specialized Gemini models in natural language (for more details see Appendix A).
@@
    Multimodal training from scratch with an additional image input - a diagram of the given geometry problem (for more details see Appendix B).
@@
All these sets contain full proofs, and during training we compute perplexity loss on them. Note, however, that these are only proxy metrics for two reasons. First, during inference (just like in AG1) we only use auxiliary points suggested by the language model, while the perplexity is computed on the entire proof. Second, there might be multiple ways to solve a given problem, but perplexity is computed for one particular solution. Just like in AG1, our main downstream metric is the solve rate on IMO problems, where the language model produces auxiliary points followed by a DDAR run via beam search described in section 6. These results will be discussed in Section 8.
We train our models with the largest possible batch size allowed by the hardware [^4] We did not observe any training issues compared to smaller batches. using TPUv4. A learning rate schedule is a linear warm-up followed by the cosine anneal. Learning rate hyperparameters are determined from scaling laws. On Figure 5 we illustrate learning curves for different size Gemini models in terms of parameter count. As expected, increasing the model size decreases perplexity loss for train, eval and our special IMO evaluation set.
Figure 5: Learning curves for AlphaGeometry2 language models of different sizes in terms of parameter count (“m" - million, “B" - billion). Increasing the model size results in decreasing loss for train, eval and IMO evaluation sets.

### v3

@@ Abstract
We present AlphaGeometry2 (AG2), a significantly improved version of AlphaGeometry introduced in Trinh2024, which has now surpassed an average gold medalist in solving Olympiad geometry problems. To achieve this, we first extend the original AlphaGeometry language to tackle problems involving movements of objects, and problems containing linear equations of angles, ratios, and distances. This, together with support for non-constructive problems, has markedly improved the coverage rate of the AlphaGeometry language on International Math Olympiads (IMO) 2000-2024 geometry problems from 66% to 88%. The search process of AG2 has also been greatly improved through the use of Gemini architecture for better language modeling, and a novel knowledge-sharing mechanism that enables effective communication between search trees. Together with further enhancements to the symbolic engine and synthetic data generation, we have significantly boosted the overall solving rate of AG to 84% on all geometry problems over the last 25 years, compared to 54% previously. AG2 was also part of the system that achieved the silver-medal standard at IMO 2024 https://deepmind.google/blog/ai-solves-imo-problems-at-silver-medal-level/. Finally, we report progress towards using AG2 as a part of a fully automated system that reliably solves geometry problems from natural language input. Code: https://github.com/google-deepmind/alphageometry2.
@@ 1 Introduction
Advancing reasoning capabilities of artificial intelligence (AI) systems, especially in tackling complex mathematical problems, has been a decade-long holy grail of AI research. Even with the rapid developments of large language models, most systems until now still struggle with the fundamental understanding of geometric objects (e.g., points, lines, circles) and their interactions. This had motivated our previous work, AlphaGeometry Trinh2024, a neuro-symbolic system capable of solving Euclidean geometry problems at the Olympiad level. Specifically, AlphaGeometry (AG1) combines the power of a neural language model, that is creative in suggesting insights for tackling challenging problems, with a symbolic engine, that uses a formal language to describe geometric constructs and deduction rules to reliably derive new properties. Trained on 100M synthetically generated examples of diverse levels of complexity, AG1 was able to solve geometry problems at the level of the International Mathematical Olympiad (IMO), a prestigious competition for the world’s brightest high-school mathematicians.
@@
Broadly speaking, there are two main approaches for automatically solving geometry problems. The first group is about bashing the geometry problems algebraically with Wu’s method chou1985proving; wu2008decision, Area method chou1993automated; chou1994machine, or Gröbner bases kapur1986geometry; kapur1986using. The second group, on the other hand, relies on synthetic techniques, such as Deduction database chou2000deductive, or Full angle method chou1996automated. AG1 uses the latter, as a more human-like approach suitable for transferring the research knowledge to other domains. The neuro-symbolic system that AG1 employed has demonstrated a significant step towards mastering the domain of Euclidean geometry. However, despite its success, AG1 exhibited limitations in several key areas. Its performance was constrained by the scope of its domain-specific language, the efficiency of the symbolic engine, and the capacity of the initial language model. As a result, when considering all the recent IMO geometry problems from the year 2000 until now, AG1 can only achieve a solving rate of 54%.
@@
This paper introduces AlphaGeometry2 (AG2), a substantial upgrade that addresses the aforementioned limitations and significantly enhances performance. First, we expand the domain language to encompass a wider range of geometric concepts, including locus theorems with moving constructs and linear equations of angles, ratios, and distances. We also introduce a significantly faster and more robust symbolic engine, incorporating optimizations such as a reduced rule set and enhanced handling of double points. AG2 leverages a more powerful language model based on Gemini team2024gemini and trained on an order of magnitude larger and more diverse dataset. To further improve performance, we develop a novel search algorithm, Shared Knowledge Ensemble of Search Trees (SKEST), that explores a broader range of auxiliary construction strategies and employs a knowledge-sharing mechanism to expand and accelerate the search process. Finally, we make progress towards building a fully automated and reliable system that solves geometry problems in natural language. To do this we utilize Gemini to translate problems from natural language into the AlphaGeometry language and implement a new automated diagram generation algorithm.
@@
  • Stronger and faster symbolic engine: introducing an optimized rule set, better handling of double points, and a two-order-of-magnitude faster implementation in C++.
@@
For related work and a discussion on how the AlphaGeometry2 positions among other systems reported in the literature, we refer the readers to Appendix A.
@@ 2 More general domain language
First introduced in Trinh2024, AG1 uses a simple domain-specific language that consists of nine basic “predicates” listed in Table 1. While these predicates are sufficient to cover 66% of all geometry problems for 2000-2024 IMO, AG1 language does not allow talking about linear equations, movements of points/lines/circles or common computational problems such as “Find the angle …” (e.g., IMO 2009 P4). Below, we explain how AG2 addresses these and other challenges.
@@
eqangle a b c d e f g h Directed angle between $AB$ and $CD$ is equal to the one between $EF$ and GHGH
@@ 3.2 Faster algorithm
The DDAR1 algorithm is processing a list of rules and tries to apply each rule to all combinations of points. This process involves a candidate searching step, whose time complexity is polynomial in the number of points, and a clause matching step, whose time complexity is exponential in the number of clauses per premise. In theory, the worst case for searching similar triangle candidates in AG1 is $O(N^{8})$, which is one of the most time-consuming steps. The exponential clause matching step is another expensive step. To make the search more efficient, we take all essential rules and hard-code search for their application, which reduces the number of queries for the AR sub-engine to at most cubic. Furthermore, we discard the explicit rules for angles and distances (e.g., about perpendicular or parallel lines) – all such deductions happen automatically in the AR engine.
@@ 3.3 Faster implementation
While the new algorithm already significantly accelerates DDAR, we make further speed improvements by implementing its core computation (Gaussian Elimination) in C++. The new C++ library, which is exported into Python via pybind11 pybind11, is over 300 times faster than DDAR1. In order to benchmark the speed improvement, we select a set of 25 IMO problems that cannot be solved by DDAR (see Figure 8), and run the test 50 times on a machine with AMD EPYC 7B13 64-core CPU. While on average DDAR1 finishes its computations in $1179.57\pm 8.055$ seconds, DDAR2 is much faster - finishing in $3.44711\pm 0.05476$ seconds [^1] The average running time may vary depending on the machine status at different times..
@@ 4 Better synthetic training data
Supplementing the symbolic engine with a language model was a key to AG1’s success, bringing the solve rate from 14 (pure deduction proofs) to 25 out of 30 selected IMO problems Trinh2024. This language model was trained on a large amount of algorithmically generated synthetic data. In AG2, we use the same procedure.
Similar to AG1, our synthetic data generation method starts by sampling a random diagram, and uses the symbolic engine to deduce all possible facts from it. For each of the deduced facts, a traceback algorithm is used to extract the corresponding premises, auxiliary points, and deduction steps that prove the fact. Our data generation method deliberately avoids the use of human-crafted problems as initial diagram seeds, and strictly starts from random diagrams. This design choice eliminates the risk of data contamination and allows for the exploration of theorem distributions that may extend beyond established human knowledge. This approach contrasts with methods like TongGeometry zhang2024proposing, which rely on human expertise and existing problem diagrams to guide and filter data generation. In AG2, we keep using random diagrams as initial seeds and continue to push for better synthetic training data.
@@
We also improved the speed of the data generation algorithm. Recall that in AG1, we first run deduction closure on random diagrams, and “traceback" to obtain the minimal problem and minimal proof that proves each fact in the closure. To obtain the minimal problem in AG1, we have to exhaustively remove different subsets of points from the problem and rerun DDAR to check provability. Such a search can find the subset with the smallest cardinality, however, as an exponential search, it is infeasible for a larger number of points. Therefore, we switched to a greedily discarding algorithm shown in Figure 3, which uses just a linear number of checks for whether a set of points suffices to prove the goal. The greedy algorithm is guaranteed to find a minimal set of points with respect to inclusion as long as the check is monotonic (if $A\subseteq B$, then $\texttt{check\_provable}(A)\Rightarrow\texttt{check\_provable}(B)$). In reality, we also require the pruned set to remain closed under construction dependencies (so that we can still run a random construction). If we incorporate this condition into the check_provable predicate, it stops being monotonic. This difficulty can be fixed by processing the points via the algorithm from Figure 3 in a reverse-topological order (first points that do not depend on any other points, and last the initial points of the construction).
@@ 5 Novel search algorithm
In AG1, we use a simple beam search to discover proofs. In AG2, we design a novel search algorithm, in which several differently configured beam searches are executed in parallel and are allowed to help each other through a knowledge sharing mechanism (see Figure 4). To improve the robustness of our system, we use multiple different language models for each search tree configuration. We call this search algorithm Shared Knowledge Ensemble of Search Trees (SKEST).
@@
  • “Classic" search tree: the same beam tree search used in AG1, where a language model is asked to produce one auxiliary point at each node.
  • Tree predicting multiple auxiliary points at each node: a language model is allowed to produce as many auxiliary points as it wants at each tree node. Recall that this is possible because our LM is trained to produce full proofs, starting with auxiliary points and followed by the deduction steps [^2] See a more detailed discussion on producing full proofs with a language model alone in Section F. Note that even though we want our models to generate all necessary auxiliary points in one query, in practice, we observe the need to call the model multiple times given previously produced auxiliary points. Allowing the model to produce multiple auxiliary points accelerates finding a solution and effectively increases the tree search depth.
  • Tree predicting different types of aux points uniformly. Recall that LM outputs for auxiliary points look like x00 a : cong a b c d (00) coll a e f (01), i.e. “construct point a such that a b = c d and a e f are collinear". Typically, to predict aux points, we prompt the language model with the first token x00 and let it generate the rest. Here, instead, we prompt the LM with x00 a : cong, x00 a : coll, x00 a : cyclic, x00 a : perp, etc. to force uniform distribution across the first 4 tokens, and then let the LM generate the rest.
@@ 6.1 Training setup
AG1 language model, a custom transformer, was trained in an unsupervised fashion in two phases: training on problems with and without auxiliary constructions followed by training on only problems that contain auxiliary constructions. For AG2, we leverage the Gemini training pipeline and simplify training to just one phase: unsupervised learning on all data. Our new language model is a sparse mixture-of-expert Transformer-based model that builds on Gemini team2024gemini and is trained on AG2 data described in Section 4. We train multiple models of different sizes using three training setups:
@@
    Training from scratch with a custom tokenizer in the domain-specific language (AG1 setup).
@@
    Fine-tuning already pre-trained custom math specialized Gemini models in natural language (for more details see Appendix B).
@@
    Multimodal training from scratch with an additional image input - a diagram of the given geometry problem (for more details see Appendix C).
@@
All these sets contain full proofs, and we compute perplexity loss on them during training. Note, however, that these are only proxy metrics for two reasons. First, during inference (just like in AG1), we only use auxiliary points suggested by the language model, while the perplexity is computed on the entire proof. Second, there might be multiple ways to solve a given problem, but perplexity is computed for one particular solution. Just like in AG1, our main downstream metric is the solve rate on IMO problems, where the language model produces auxiliary points followed by a DDAR run via beam search described in section 5. These results will be discussed in Section 7.
We train our models with the largest possible batch size allowed by the hardware [^4] We did not observe any training issues compared to smaller batches. using TPUv4. A learning rate schedule is a linear warm-up followed by the cosine anneal. Learning rate hyperparameters are determined from scaling laws. On Figure 5, we illustrate learning curves for different-sized Gemini models in terms of parameter count. As expected, increasing the model size decreases perplexity loss for train, eval and our special IMO evaluation set.
Figure 5: Learning curves for AlphaGeometry2 language models of different sizes in terms of parameter count (“m" - million, “B" - billion, for example, “3p3B” means a model with 3.3 billion parameters). Increasing the model size results in decreasing loss for train, eval and IMO evaluation sets.